Five times the sum of 3 and twice some number is 70. What is the number?
Found 2 solutions by CharlesG2, solver91311:
Answer by CharlesG2(834)

Five times the sum of 3 and twice some number is 70. What is the number?

let x = some number
sum of 3 and twice some number = 3 + 2x
five times the sum of 3 and twice some number = 5 * (3 + 2x)


5 * (3 + 2x) = 70
15 + 10x = 70 (distributed the 5)
10x = 55 (subtracted 15 from both sides)
x = 55/10 = 5.5


Answer by solver91311(24713)



Let represent the "some number." Then twice some number must be . Then the sum of 3 and twice some number must be . Then 5 times the sum...must be . Finally, "is" means equals so:



Solve for .

John

My calculator said it, I believe it, that settles it